OpenStudy (zepp):

Hello! What is the easiest way to study the convergence or divergence of this serie? \[\sum _{n=1}^{\infty \:}\left(1+\frac{1}{n}\right)^{n^2}\]

OpenStudy (jango_in_dtown):

@zepp Cauchy's root test

OpenStudy (jango_in_dtown):

\[(1+1/n)^{n^2}=((1+1/n)^{n})^{n}\]

OpenStudy (zepp):

Oh! I didn't see that, I'll work from there, thanks!

OpenStudy (jango_in_dtown):

Hints, the limit is e . now apply ratio test. and tell me the answer

OpenStudy (jango_in_dtown):

e>1

OpenStudy (zepp):

Yeah, by the root test test, \(\lim_{x \rightarrow \infty}\ |{a_{n}^{\frac{1}{n}}}|=e>1\) thus the series diverges

OpenStudy (jango_in_dtown):

right

OpenStudy (zepp):

Thank you!